# Variability of X-ray polarization of Cyg X-1

### v1

@@
We present the results of a three-year X-ray, optical, and radio polarimetric monitoring campaign of the prototypical black hole X-ray binary Cyg X-1, conducted from 2022 to 2024. The X-ray polarization of Cyg X-1 was measured 13 times with the Imaging X-ray Polarimetry Explorer (IXPE), covering both hard and soft spectral states. The X-ray polarization degree (PD) in the hard state was found to be 4.0%, roughly twice as high as in the soft state, where it was around 2.2%. In both states, a statistically significant increase of PD with the energy was found. Moreover, a linear relation between PD and spectral hardness suggests a gradual and continuous evolution of the polarization properties, rather than an abrupt change of polarization production mechanism between states. The polarization angle (PA) was independent of the spectral state and showed no trend with the photon energy. The X-ray PA is well aligned with the orientation of the radio jet, as well as the optical and radio PAs. We find significant orbital changes of PA in the hard state, which we attribute to scattering of X-ray emission at intrabinary structure. No significant superorbital variability in PD or PA was found at the period d. We also find no correlation between the X-ray and optical polarization; if any, there is a long-term anti-correlation between the X-ray PD and the radio PD.
@@
Accretion is an efficient mechanism for heating and extracting energy from matter as it falls toward a compact object. Energy can be released through various channels, producing distinct X-ray emission signatures (spectral states) observed in these systems. Black hole X-ray binaries (BHXRBs) are known to swing between two major spectral states, hard and soft, distinguished by the energies at which their emission peaks (Zdziarski & Gierliński, 2004). The soft state is characterized by dominant thermal emission from the classical optically thick and geometrically thin accretion disk (Shakura & Sunyaev, 1973; Novikov & Thorne, 1973). Transition to the hard spectral state is marked by spectral shift to a power-law shape, peaking at keV. This emission is believed to be produced by multiple Compton upscatterings of soft seed photons, either from the disk or internally-produced synchrotron photons, in a hot optically thin medium surrounding the black hole (Sunyaev & Titarchuk, 1980; Gierliński et al., 1997; Poutanen & Vurm, 2009; Veledina et al., 2013). Determining the size of the hot, rarefied plasma responsible for the bulk of X-ray emission, along with its shape and orientation with respect to the disk, remains one of the key open questions in high-energy astrophysics.
@@
The detected PD in Cyg X-1 is however too high for the known system inclination, and is hard to achieve in any scenario. Additional assumptions have been employed to achieve PD% in the models: the inclination of the X-ray emitting region might be higher (by –) than the orbital one (Krawczynski et al., 2022), the hot medium may attain a significant outflow velocity (Poutanen et al., 2023), or the PD produced intrinsically in the source may be boosted by the scattering off the accretion disk wind at large radii (Nitindala et al., 2025).
@@ 2.1 IXPE
We consider the full set of IXPE observations of Cyg X-1, which consists of 13 individual pointings, conducted from 2022 to 2024 (see Table 1). We have used Level 2 data downloaded from the IXPE archive at HEASARC and analyzed it using ixpeobssim package v31.0.1 (Baldini et al., 2022). In the case of the hard state observation in May 2022 we have used post-reconstruction calibration of the energy scale.111https://heasarc.gsfc.nasa.gov/FTP/ixpe/data/obs/01/01002901/README The source region for all available observations was defined in SAOimageDS9 v8.6 as a circle with radius around the source and extracted using xpselect tool. The average polarization properties in the entire energy band for each observation were extracted using the xpbin tool with PCUBE algorithm in a single 2–8 keV energy bin. For phase-resolved polarimetric analysis, we utilized the xpphase tool with the orbital period of 5.599829 d and a zero point of JD 2441874.707 (Brocksopp et al., 1999b). We split Epoch 1 data into 10 phase bins with xpselect and calculated the polarization properties in each phase bin with the same approach as for the whole observation. The hardness ratio was calculated as IXPE photon flux ratio in the energy ranges 4–8 and 2–4 keV. The main results are shown in Fig. 1.
Spectra were extracted with the xpbin PHA1 algorithm, utilizing CalDB v13 response files. The spectra were then fitted with the X-ray fitting package xspec v12.14.1 (Arnaud, 1996), using a simple model constant(diskbb+powerlaw). The superorbital phases were calculated with a period of 294.0 d and a zero point of JD 2440000.0 (Priedhorsky et al., 1983).
@@ 2.2.1 Optical observations
Optical polarimetric observations of Cyg X-1 were carried out with the broad-band polarimeters DIPol-2 (Piirola et al., 2014) and DIPol-UF (Piirola et al., 2021), mounted on the remotely-controlled 60 cm Tohoku telescope (T60) at Haleakala Observatory, Hawaii and at the 2.56 m Nordic Optical Telescope (NOT), Observatorio del Roque de los Muchachos (ORM), La Palma, Spain, and with RoboPol polarimeter (Ramaprakash et al., 2019) in the focal plane of the 1.3 m telescope of the Skinakas observatory, Greece. DIPol-2 and DIPol-UF are high-precision double-image CCD polarimeters, capable of measuring polarization simultaneously in three optical () bands. The instrumental polarization of both instruments is small () and is well calibrated by observing 15 to 20 unpolarized standard stars. The polarization of the sky is optically eliminated by the design of the instruments. The zero point of the PA was determined by observing highly polarized standards HD 204827 and HD 161056. Each measurement of Stokes parameters took about 20 s and more than 200 individual measurements were obtained during the average observing night. In total, Cyg X-1 was observed for 90 nights in 2022–2024. Intrinsic polarization of the source has been extracted by subtracting the Stokes parameters of the interstellar polarization (Table 2 of Kravtsov et al., 2023) from the observed Stokes parameters of Cyg X-1. A more detailed description of the methods and calibrations can be found in Piirola et al. (2020) and Kravtsov et al. (2023). The RoboPol data was analyzed by the automatic pipeline described in Blinov et al. (2021) using both polarized and unpolarized standards to characterize the instrumental polarization. For the sources not in the central mask, used to account for the ISM polarization, we use the analysis procedure described in Panopoulou et al. (2015). For consistency, the ISM polarization is corrected using the same reference star for all three instruments.
@@ 2.2.3 RATAN 600
We carried out monitoring of Cyg X-1 with the RATAN 600 radio telescope at 4.7 GHz and 11.2 GHz from 2022 May to 2024 June using the ‘Southern Sector and Flat mirror’ antenna. The sensitivity of such measurements is about 3–10 mJy beam-1. Thus, Cyg X-1 was undetected most of the time, with detections presented in Fig. 2. Previous monitoring observations of Cyg X-1 have shown typical flux variations in the vicinity of 10–30 mJy at 4.7 GHz. Calibration was performed using quasar 3C 48, adopting a brightness of 5.8 and 3.42 Jy at 4.7 and 8.2 GHz, respectively, according to the flux density scale by Ott et al. (1994).
@@ 2.2.4 AMI
Cyg X-1 was observed 98 times during May and June 2023 and May and June 2024 with the AMI Large Array (Zwart et al., 2008; Hickish et al., 2018) at 15.5 GHz. The observations were typically minutes, with two 10 minute scans of Cyg X-1 interleaved between short observations of a nearby compact source. The flux density scale of the observations was set by using daily short observations of 3C 286, and the interleaved calibrator observations were used to calibrate antenna-based amplitude and phase variations during the observations. The observations covered a 5 GHz bandwidth of a single linear polarization, Stokes .
@@ 3.1 Long-term X-ray polarization changes
The energy dependence of polarization, averaged separately over the hard and soft states, is shown in Fig. 3. The PD grows with energy in both states, as indicated by the highly statistically significant improvement of the fit with linearly increasing PD with energy comparing to the constant PD model: the /d.o.f. value drops from 27.5/5 to 6.3/4 in the hard state and from 78/5 to 7/4 in the soft state.
We further investigate the dependence of PD and PA on hardness in Fig. 4. We find that the X-ray PD shows strong positive correlation (Pearson correlation coefficient ) with spectral hardness, while PA is not sensitive to spectral changes. We considered two alternative scenarios for the PD behavior with hardness: two constant levels of polarization, that can be motivated by the change of the dominant mechanism responsible for polarization in the hard and soft state (e.g., returning radiation in the soft state and Comptonization in the hard state), as opposed to linear trend with hardness, which may correspond to a gradual changes of parameters of the flow (e.g., if Comptonization could be responsible for the polarization in both states). For the two constant levels we get /d.o.f.=21/11 and for the linear trend we get /d.o.f.=13.6/11, which is preferred with high significance.
@@ 3.2 Orbital variability
The orbital motion of an X-ray source in the presence of a companion star and circumstellar (intrabinary) matter can induce variations of the observed polarization on the orbital timescale (Brown et al., 1978; Kravtsov et al., 2020; Rankin et al., 2024; Ahlberg et al., 2024). To verify the presence of variability at the orbital period in IXPE observations of Cyg X-1, we performed a phase-resolved polarimetric analysis. The dependence of PD on the spectral hardness does not allow us to combine the data in all spectral states. To minimize the influence of spectral hardness on polarization, we treated soft- and hard-state data separately.
We primarily focus on the hard-state data owing to repeating coverage of the same orbital phases over years and one complete orbital cycle (Epoch 1). We split the Epoch 1 data into 10 phase bins as described in Sect.2.1, while the other, shorter observations were not subdivided. The results are shown in Figs. 5 and 6, where variations of the Stokes and are presented as function of the orbital phase and on the -plane, and also through variation of PD and PA. The X-ray polarization of Cyg X-1 shows hints of orbital variability: the PD varies sinusoidally, gradually increasing from 3% at orbital phase 0.1 to the maximum of 5% at phase 0.5. The observed PA displays a swing with an amplitude of 10°, roughly half-cycle out of phase with the PD. This behavior can be interpreted as the tip of the polarization vector tracing a closed loop in the -plane (see Fig. 5, left panel).
@@ 3.2.1 Constant versus sine-wave
To assess the statistical significance of the orbital variations, we used the F-test to compare two nested models. The null hypothesis, , assumes that the Stokes parameters and remain constant across all orbital phases. It has 2 free parameters: and . We find the /d.o.f. = 30.5/28 for the fit with this model (see black dashed line in Fig. 6).
@@
where indices (c) and (v) denote constant and variable components, respectively, and is the zero-phase angle. The model has four free parameters: , , , and . In this model, the trajectory the points track at the -plane may be decomposed into the sum of two vectors: the first one, constant vector, links the origin with the average polarization (center of the circle), and another, varying vector, with constant length of and with the azimuthal angle being related to the orbital phase , links the center of the circle with the instantaneous parameters. Fig. 6 shows the best-fit model with the solid orange lines that gives /d.o.f.=24.2/26.
@@ 3.2.2 Rotating vector model
From the observed normalized Stokes parameters we subtract the average values to obtain a new set of Stokes parameters representing the variable component:
We calculate and from the Epoch 1 data only (10 blue data points in Fig. 5), because it continuously covers the whole orbital period – this ensures that the center point of the loop will not be biased due to unequal orbital coverage. We then compute PDv and PAv of the variable component using standard formulae. The phase-resolved variations of PAv are shown in Fig. 7. We see that PAv shows pronounced, gradual rotation along the orbit (corresponding to one full loop in the left panel of Fig. 5). Here, for the sake of an easier interpretation of the observed results, we show , which is the difference between PAv and the PA of the constant component (obtained from and ). One can think of as the deviations of the polarization orientation of the variable component from the average direction of polarization on the sky (i.e., relative to the position angle of the orbital axis, rather than relative to the direction towards North).
@@ 3.3 Superorbital variability
Next, we test presence of super-orbital variability in the X-ray polarization data. The period of these modulations was reported to vary over time: both 300 d and 150 d were found (e.g., Brocksopp et al., 1999a; Lachowicz et al., 2006; Ibragimov et al., 2007; Zdziarski et al., 2011). The nature of these variations remains uncertain, but the scenario with precession of the accretion disk or its inner parts is consistent with variations of several observables simultaneously (Bochkarev & Karitskaya, 1983; Poutanen et al., 2008). If true, this model predicts coordinated variations of the X-ray polarimetric signatures. These variations, however, can be mixed with the orbital variability, hence polarization from different superorbital phases should be compared for the same orbital bins. Moreover, because of the change of the average PD between hard and soft states, polarization changes cannot be considered jointly for both states.
To test the precessing inner flow scenario, we fold the hard-state X-ray polarization measurements according to the phase of the superorbital period d with zero point JD 2440000 (Priedhorsky et al., 1983; Kemp et al., 1983).444No significant variations of X-ray polarization were previously found over d (Krawczynski et al., 2022), which corresponds to the strongest long-term period observed in MAXI (Matsuoka et al., 2009) light curves, but had an instrumental origin. In Fig. 5 we show the resulting PD and PA color-coded according to the superorbital phase. Current data show no obvious dependence of the polarization properties on the superorbital phase: for the same orbital phase bins, both PD and PA are consistent with being constant for different superorbital phases. The only exception is the PA of Epoch 11 (orange cross centered at orbital phase 0.9), which however has similar super-orbital phase to the Epochs 9 and 10 (centered at orbital phases 0.9 and 0.4, respectively), which are well aligned with other points at different superorbital phases (blue crosses). This indicates that the PA of the outlier Epoch 11 is not directly related to the superorbital changes. Current data suggest that, if present, the superorbital changes of PA do not exceed with confidence. This value is substantially smaller than the previously-suggested misalignment angle between the orbital and inner flow axes 15°–20° needed to explain the observed variations of fluxes and polarization signatures.

### v2

@@
We present the results of a three-year X-ray, optical, and radio polarimetric monitoring campaign of the prototypical black hole X-ray binary Cyg X-1, conducted from 2022 to 2024. The X-ray polarization of Cyg X-1 was measured 13 times with the Imaging X-ray Polarimetry Explorer (IXPE), covering both hard and soft spectral states. The X-ray polarization degree (PD) in the hard state was found to be 4.0%, roughly twice as high as in the soft state, where it was around 2.2%. In both states, a statistically significant increase in PD with the energy was found. Moreover, a linear relation between PD and spectral hardness suggests a gradual and continuous evolution of the polarization properties, rather than an abrupt change of polarization production mechanism between states. The polarization angle (PA) was independent of the spectral state and showed no trend with the photon energy. The X-ray PA is well aligned with the orientation of the radio jet, as well as the optical and radio PAs. We find significant orbital changes of PA in the hard state, which we attribute to scattering of X-ray emission at the intrabinary structure. No significant superorbital variability in PD or PA was found at the period d. We detect, for the first time in this source, polarization of the radio emission, with the PA aligned with the jet, and a strong increase of the PD at a transition to the soft state. We also find no correlation between the X-ray and optical polarization; if any, there is a long-term anti-correlation between the X-ray PD and the radio PD.
@@
Accretion is an efficient mechanism for heating and extracting energy from matter as it falls toward a compact object. Energy can be released through various channels, producing distinct X-ray emission signatures (spectral states) observed in these systems. Black hole X-ray binaries (BHXRBs) are known to swing between two major spectral states, hard and soft, distinguished by the energies at which their emission peaks (Zdziarski & Gierliński, 2004). The soft state is characterized by dominant thermal emission from the classical optically thick and geometrically thin accretion disk (Shakura & Sunyaev, 1973; Novikov & Thorne, 1973). Transition to the hard spectral state is marked by spectral shift to a power-law shape, peaking at keV. This emission is believed to be produced by multiple Compton upscatterings of soft seed photons, either from the disk or internally produced synchrotron photons, in a hot optically thin medium surrounding the black hole (Sunyaev & Titarchuk, 1980; Gierliński et al., 1997; Poutanen & Vurm, 2009; Veledina et al., 2013). Determining the size of the hot, rarefied plasma responsible for the bulk of X-ray emission, along with its shape and orientation with respect to the disk, remains one of the key open questions in high-energy astrophysics.
@@
The detected PD in Cyg X-1 is, however, too high for the known system inclination, and is hard to achieve in any scenario. Additional assumptions have been employed to achieve PD% in the models: the inclination of the X-ray emitting region might be higher (by –) than the orbital one (Krawczynski et al., 2022), the hot medium may attain a significant outflow velocity (Poutanen et al., 2023), or the PD produced intrinsically in the source may be boosted by the scattering off the accretion disk wind at large radii (Nitindala et al., 2025).
@@ 2.1 IXPE
We consider the full set of IXPE observations of Cyg X-1, which consists of 13 individual pointings, conducted from 2022 to 2024 (see Table 1). We used Level 2 data downloaded from the IXPE archive at HEASARC and analyzed it using ixpeobssim package v31.0.1 (Baldini et al., 2022). In the case of the hard state observation in May 2022 we used post-reconstruction calibration of the energy scale.111https://heasarc.gsfc.nasa.gov/FTP/ixpe/data/obs/01/01002901/README The source region for all available observations was defined in SAOimageDS9 v8.6 as a circle with radius around the source and extracted using xpselect tool. The average polarization properties in the entire energy band for each observation were extracted using the xpbin tool with PCUBE algorithm in a single 2–8 keV energy bin. For phase-resolved polarimetric analysis, we utilized the xpphase tool with the orbital period of 5.599829 d and a zero point of JD 2441874.707 (Brocksopp et al., 1999b). We split Epoch 1 data into ten phase bins with xpselect and calculated the polarization properties in each phase bin with the same approach as for the whole observation. The hardness ratio was calculated as IXPE photon flux ratio in the energy ranges 4–8 and 2–4 keV. The main results are shown in Fig. 1. Spectra were extracted with the xpbin PHA1 algorithm, utilizing CalDB v13 response files. The superorbital phases were calculated with a period of 294.0 d and a zero point of JD 2440000.0 (Priedhorsky et al., 1983).
@@ 2.2.1 Optical observations
Optical polarimetric observations of Cyg X-1 were carried out with the broad-band polarimeters DIPol-2 (Piirola et al., 2014) and DIPol-UF (Piirola et al., 2021), mounted on the remotely controlled 60 cm Tohoku telescope (T60) at Haleakala Observatory, Hawaii and at the 2.56 m Nordic Optical Telescope (NOT), Observatorio del Roque de los Muchachos (ORM), La Palma, Spain, and with RoboPol polarimeter (Ramaprakash et al., 2019) in the focal plane of the 1.3 m telescope of the Skinakas observatory, Greece. DIPol-2 and DIPol-UF are high-precision double-image CCD polarimeters, capable of measuring polarization simultaneously in three optical () bands. The instrumental polarization of both instruments is small () and is well calibrated by observing 15 to 20 unpolarized standard stars. The polarization of the sky is optically eliminated by the design of the instruments. The zero point of the PA was determined by observing highly polarized standards HD 204827 and HD 161056. Each measurement of Stokes parameters took about 20 s and more than 200 individual measurements were obtained during the average observing night. In total, Cyg X-1 was observed for 90 nights in 2022–2024. Intrinsic polarization of the source has been extracted by subtracting the Stokes parameters of the interstellar polarization (Table 2 of Kravtsov et al., 2023) from the observed Stokes parameters of Cyg X-1. A more detailed description of the methods and calibrations can be found in Piirola et al. (2020) and Kravtsov et al. (2023). The RoboPol data was analyzed by the automatic pipeline described in Blinov et al. (2021) using both polarized and unpolarized standards to characterize the instrumental polarization. For the sources not in the central mask, used to account for the ISM polarization, we used the analysis procedure described in Panopoulou et al. (2015). For consistency, the ISM polarization is corrected using the same reference star for all three instruments.
@@ 2.2.3 RATAN 600
We carried out monitoring of Cyg X-1 with the RATAN 600 radio telescope at 4.7 GHz and 11.2 GHz from 2022 May to 2024 June using the ”Southern Sector and Flat mirror” antenna. The sensitivity of such measurements is about 3–10 mJy beam-1. Thus, Cyg X-1 was undetected most of the time, with detections presented in Fig. 2. Previous monitoring observations of Cyg X-1 have shown typical flux variations in the vicinity of 10–30 mJy at 4.7 GHz. Calibration was performed using quasar 3C 48, adopting a brightness of 5.8 and 3.42 Jy at 4.7 and 8.2 GHz, respectively, according to the flux density scale by Ott et al. (1994).
@@ 2.2.4 AMI
Cyg X-1 was observed 98 times during May and June 2023 and May and June 2024 with the AMI Large Array (Zwart et al., 2008; Hickish et al., 2018) at 15.5 GHz. The observations were typically minutes, with two ten-minute scans of Cyg X-1 interleaved between short observations of a nearby compact source. The flux density scale of the observations was set by using daily short observations of 3C 286, and the interleaved calibrator observations were used to calibrate antenna-based amplitude and phase variations during the observations. The observations covered a 5 GHz bandwidth of a single linear polarization, Stokes .
@@ 3.1 Long-term X-ray polarization changes
The energy dependence of polarization, averaged separately over the hard and soft states, is shown in Fig. 3. The PD grows with energy in both states, as indicated by the highly statistically significant improvement of the fit with linearly increasing PD with energy comparing to the constant PD model: the /d.o.f. value drops from 27.5/5 to 6.3/4 in the hard state and from 78/5 to 7/4 in the soft state. We further investigate the dependence of PD and PA on hardness in Fig. 4. We find that the X-ray PD shows strong positive correlation (Pearson correlation coefficient ) with spectral hardness, while PA is not sensitive to spectral changes – fit of the linear model to the PA gives /d.o.f.=13.6/11.
@@ 3.2 Orbital variability
The orbital motion of an X-ray source in the presence of a companion star and circumstellar (intrabinary) matter can induce variations of the observed polarization on the orbital timescale (Brown et al., 1978; Kravtsov et al., 2020; Rankin et al., 2024; Ahlberg et al., 2024). To verify the presence of variability at the orbital period in IXPE observations of Cyg X-1, we performed a phase-resolved polarimetric analysis. The dependence of PD on the spectral hardness prevents us from combining the data in all spectral states. To minimize the influence of spectral hardness on polarization, we treated soft- and hard-state data separately.
We primarily focus on the hard-state data owing to repeating coverage of the same orbital phases over years and one complete orbital cycle (Epoch 1). We split the Epoch 1 data into 10 phase bins as described in Sect.2.1, while the other, shorter observations were not subdivided. The results are shown in Figs. 5 and 6, where variations of the Stokes and are presented as function of the orbital phase and on the -plane, and also through variation in PD and PA. The X-ray polarization of Cyg X-1 shows hints of orbital variability: the PD varies sinusoidally, gradually increasing from 3% at orbital phase 0.1 to the maximum of 5% at phase 0.5. The observed PA displays a swing with an amplitude of 10°, roughly half-cycle out of phase with the PD. This behavior can be interpreted as the tip of the polarization vector tracing a closed loop in the -plane (see Fig. 5, left panel).
@@ 3.2.1 Constant versus sine-wave
To assess the statistical significance of the orbital variations, we used the F-test to compare two nested models. The null hypothesis, , assumes that the Stokes parameters and remain constant across all orbital phases. It has two free parameters: and . We find the /d.o.f. = 30.5/28 for the fit with this model (see dashed black line in Fig. 6).
@@
where indices (c) and (v) denote constant and variable components, respectively, and is the zero-phase angle. The model has four free parameters: , , , and . In this model, the trajectory the points track at the -plane may be decomposed into the sum of two vectors: the first one, constant vector, links the origin with the average polarization (center of the circle), and another, varying vector, with constant length of and with the azimuthal angle being related to the orbital phase , links the center of the circle with the instantaneous parameters. Fig. 6 shows the best-fit model with the solid blue lines that gives /d.o.f.=24.2/26.
@@ 3.2.2 Rotating vector model
From the observed normalized Stokes parameters we subtracted the average values to obtain a new set of Stokes parameters representing the variable component:
We calculated and from the Epoch 1 data only (ten blue data points in Fig. 5), because it continuously covers the whole orbital period – this ensures that the center point of the loop will not be biased due to unequal orbital coverage. We then computed PDv and PAv of the variable component using standard formulae
The phase-resolved variations of PAv are shown in Fig. 7. We see that PAv shows pronounced, gradual rotation along the orbit (corresponding to one full loop in the left panel of Fig. 5). Here, for the sake of an easier interpretation of the observed results, we show , which is the difference between PAv and the of the constant component (obtained from and using Eq. 3)
One can think of as the deviations of the polarization orientation of the variable component from the average direction of polarization on the sky (i.e., relative to the position angle of the orbital axis, rather than relative to North).
@@ 3.3 Superorbital variability
Next, we tested presence of super-orbital variability in the X-ray polarization data. The period of these modulations was reported to vary over time: both 300 d and 150 d were found (e.g., Brocksopp et al., 1999a; Lachowicz et al., 2006; Ibragimov et al., 2007; Zdziarski et al., 2011). The nature of these variations remains uncertain, but the scenario with precession of the accretion disk or its inner parts is consistent with variations of several observables simultaneously (Bochkarev & Karitskaya, 1983; Poutanen et al., 2008). If true, this model predicts coordinated variations of the X-ray polarimetric signatures. These variations, however, can be mixed with the orbital variability, hence polarization from different superorbital phases should be compared for the same orbital bins. Moreover, because of the change of the average PD between hard and soft states, polarization changes cannot be considered jointly for both states.
To test the precessing inner flow scenario, we folded the hard-state X-ray polarization measurements according to the phase of the superorbital period d with zero point JD 2440000 (Priedhorsky et al., 1983; Kemp et al., 1983).444No significant variations of X-ray polarization were previously found over d (Krawczynski et al., 2022), which corresponds to the strongest long-term period observed in MAXI (Matsuoka et al., 2009) light curves, but had an instrumental origin. In Fig. 5 we show the resulting PD and PA color-coded according to the superorbital phase. Current data show no obvious dependence of the polarization properties on the superorbital phase: for the same orbital phase bins, both PD and PA are consistent with being constant for different superorbital phases. The only exception is the PA of Epoch 11 (orange cross centered at orbital phase 0.9), which, however, has similar super-orbital phase to the Epochs 9 and 10 (centered at orbital phases 0.9 and 0.4, respectively), which are well aligned with other points at different superorbital phases (blue crosses). This indicates that the PA of the outlier Epoch 11 is not directly related to the superorbital changes. Current data suggest that, if present, the superorbital changes of PA do not exceed with confidence. This value is substantially smaller than the previously suggested misalignment angle between the orbital and inner flow axes 15°–20° needed to explain the observed variations of fluxes and polarization signatures.